# Motivic classes of stacks in finite characteristic and applications to stacks of Higgs bundles

### v1

@@ 5.1. Moduli stacks of vector bundles and HN-filtrations
    There is an open substack $\mathcal{B}un_{r,d}^{\geq\tau}\subset\mathcal{B}un_{r,d}$ classifying vector bundles whose HN-type ($\tau_{1}>\ldots>\tau_{t}$) satisfies $\tau_{t}\geq\tau$;
@@
    A vector bundle $E\in\mathcal{B}un_{r,d}(K)$ is in $\mathcal{B}un_{r,d}^{\geq\tau}(K)$ if and only if there is no surjective morphism of vector bundles $E\to F$ such that the slope of $F$ is less than $\tau$;
@@ 5.2. Moduli stacks of Higgs bundles
    A Higgs bundle on $X$ is a pair $(E,\Phi)$ where $E$ is a vector bundle on $X$ and $\Phi\colon E\to E\otimes\Omega_{X}$ is an $\mathcal{O}_{X}$-linear morphism from $E$ to $E$ “twisted” by the sheaf of differential $1$-forms $\Omega_{X}$;
@@
This is an open condition compatible with field extensions, so we denote by $\mathcal{H}iggs^{ss}_{r,d}$ the open substack of semistable Higgs bundles. Furthermore, the stack $\mathcal{H}iggs^{ss}_{r,d}$ is of finite type by [FSS18, Lemma 3.3.2(ii)].
@@ 5.3. Motivic Hall algebras
In this section, we keep the assumptions from the previous section, in particular: $X$ is a smooth projective geometrically connected curve of genus $g$ over $k$. In addition we assume that there is a divisor $D$ on $X$ defined over $k$ such that $\deg D=1$.
Let $\mathcal{C}oh_{r,d}$ be the moduli stack of coherent sheaves on $X$ with rank $r$ and degree $d$. This stack is a stack locally of finite type over $k$ by [LMB00, Théorème 4.6.2.1]. For any stack $\mathcal{X}$, we consider the ring
@@
Set
@@
Let us call the group $\overline{\operatorname{Mot}}(\mathcal{X})$ defined in [FSS18, Section 2.1] without surjective and universally injective relation $\overline{\operatorname{Mot}}^{naive}(\mathcal{X})$ and the corresponding motivic Hall algebra in [FSS18, Section 5.2] $\mathcal{H}^{naive}$. Then we have the following well-defined homomorphisms
@@
Applying these homomorphisms to the identities in [FSS18, Section 5], we get identities in the “corrected” motivic Hall algebra such as [FSS18, Lemma 5.6.1 and Proposition 5.5.3]. The identity of [FSS18, Proposition 5.7.1] is also true in our $\mathcal{H}$: the argument goes through in arbitrary characteristic except for the part where the power structure is used, and this is where our Theorem 4.6 should be applied. After this, the calculations of [FSS18, Section 6] carry over to the case of finite characteristic. Thus for the rest sections, we can use the results in loc. cit.
5.4. Motivic classes of vector bundles with nilpotent endomorphisms
The following definitions come from [FSS18, Section 1.3]. We define the “normalized” zeta-function $\tilde{\zeta}_{X}(z)\coloneqq z^{1-g}\zeta_{X}(z)$ where $g$ is the genus of the curve $X$.
Recall the definition of the motivic $L$-function in Section 1.4, we define the “regularized” zeta-function by setting
$\zeta^{\ast}_{X}(\mathbb{L}^{-u}z^{v})=\begin{cases}\zeta_{X}(\mathbb{L}^{-u}z^{v})\ &\text{if}\ v>0\ \text{or}\ u>1;\\ \dfrac{L^{mot}_{X}(\mathbb{L}^{-1})}{1-\mathbb{L}^{-1}}\ &\text{if}\ (u,v)=(1,0);\\ \dfrac{L^{mot}_{X}(1)}{1-\mathbb{L}}\ &\text{if}\ (u,v)=(0,0).\end{cases}$
Let $\lambda=(\lambda_{1}\geq\lambda_{2}\geq\ldots\geq\lambda_{l}>0)$ be a partition. This can also be written as $\lambda=1^{r_{1}}2^{r_{2}}\ldots t^{r_{t}}$, where $r_{i}$ denotes the number of occurrences of $i$ among $\lambda_{j}$, $1\leq j\leq l$. The Young diagram of $\lambda$ is the set of the points $(i,j)\in\mathbb{Z}^{2}$ such that $1\leq i\leq\lambda_{j}$. For a box $s\in\lambda$, the arm $a(s)$ (resp. leg $l(s)$) is the number of boxes lying strictly to the right of (resp. strictly above) $s$.
For a partition $\lambda$, set
$J^{mot}_{\lambda}(z)=\prod_{s\in\lambda}\zeta^{\ast}_{X}(\mathbb{L}^{-1-l(s)}z^{a(s)})\in\overline{\operatorname{Mot}}(k)[[z]],$
where the product is over all boxes of the Young diagram corresponding to the partition. In particular for the empty Young diagram $\lambda$ we have $J^{mot}_{\lambda}(z)=1$.
@@ 5.4. Motivic classes of vector bundles with nilpotent endomorphisms
$L^{mot}(z_{n},\ldots,z_{1})=\frac{1}{\prod_{i<j}\tilde{\zeta}_{X}\left(\frac{z_{i}}{z_{j}}\right)}\sum_{\sigma\in S_{n}}\sigma\left\{\prod_{i<j}\tilde{\zeta}_{X}\left(\frac{z_{i}}{z_{j}}\right)\frac{1}{\prod_{i<n}\left(1-\mathbb{L}\frac{z_{i+1}}{z_{i}}\right)}\cdot\frac{1}{1-z_{1}}\right\}.$
@@
Here the sum is over all partitions $\lambda=(\lambda_{1}\geq\lambda_{2}\geq\ldots\geq\lambda_{l}>0)$, $\langle\lambda,\lambda\rangle\coloneqq\sum_{i}(\lambda^{\prime}_{i})^{2}$ where $\lambda^{\prime}=(\lambda^{\prime}_{1}\geq\ldots\geq\lambda^{\prime}_{i}\geq\ldots)$ is the conjugate partition and $|\lambda|=\sum_{i}\lambda_{i}$.
Let $\mathcal{E}nd_{r,d}^{nilp,+}$ denote the stack of HN-nonnegative bundles of rank $r$ and degree $d$ with nilpotent endomorphisms.
@@
5.5. Explicit formulas for the motivic classes of Higgs bundles
Finally we obtain the same formula for the motivic classes of semistable Higgs bundles as in [FSS18, Theorem 1.3.3]. We remove the restriction on the characteristic of the field.
@@ 5.5. Explicit formulas for the motivic classes of Higgs bundles
(5.2) $\sum_{r,d\in\mathbb{Z}_{\geq 0}\atop(r,d)\neq(0,0)}B_{r,d}w^{r}z^{d}=\mathbb{L}\operatorname{Log}\Omega_{X}^{Sch,mot}.$
@@
Recall that we have fixed a smooth projective geometrically connected curve $X$ of genus $g$. We have a result similar to [FSS18, Theorem 1.3.3] for the motivic classes of moduli stacks of semistable Higgs bundles.
@@
For the field $k$ of arbitrary characteristic, we have
@@
whenever $d$ is large enough. This formula gives an explicit answer for the volumes $|\mathcal{H}iggs^{ss}_{r,d}(\mathbb{F}_{q})|$. It is not hard to derive this formula from our Theorems 3.3 and 5.6. We suppress the details because in the next section we will give a simpler formula for $|\mathcal{H}iggs^{ss}_{r,d}(\mathbb{F}_{q})|$ based on results of Mellit.
@@ 6. Mellit’s simplification
We would like to use Mellit’s results [Mel18, Mel20a] to obtain a formula for the motivic classes of Higgs bundles in the universal $\lambda$-ring quotient that are much simpler than the formulas in Section 5.5. We also give a simpler formula for $|\mathcal{H}iggs^{ss}_{r,d}(\mathbb{F}_{q})|$.
6.1. Motivic classes of stacks of Higgs bundles in the universal $\lambda$-ring quotient
Recall from Theorem 2.18 that we have a pre-$\lambda$ structure on the ring $\overline{\operatorname{Mot}}(k)$ such that for a quasi-projective scheme $X$ we have
@@ 6.1. Motivic classes of stacks of Higgs bundles in the universal $\lambda$-ring quotient
Now we have the following proposition as a no-puncture version of [FSS24, Corollary 8.7.4].
@@ 6.2. Counting Higgs bundles over finite fields
For simplicity, we set $\alpha_{i+g}\coloneqq Q\alpha_{i}^{-1}$ for $i=1,\ldots,g$. We define a universal $L$-function
@@
$L_{X}(z)\coloneqq\left(\sum_{n\geq 0}|Sym^{n}X(\mathbb{F}_{q})|\cdot z^{n}\right)(1-z)(1-qz).$
@@
Let $d/r=d_{0}/r_{0}$ where $r_{0}$ and $d_{0}$ are coprime integers with $r_{0}>0$. The volume $|\mathcal{H}iggs^{ss}_{r,d}(\mathbb{F}_{q})|$ is the image under $ev_{X}$ of the coefficient at $w^{r}$ in
@@
Before giving the proof, we need some preliminaries. We define $V_{g}$ as the localization of $R_{g}[z]$ with respect to the multiplicative set generated by $Q^{i}-z^{j}$, where $i,j\geq 0$, $(i,j)\neq(0,0)$. Let $\mu\colon V_{g}\to\overline{\operatorname{Mot}}^{\lambda-ring}(\mathbb{F}_{q})[[z]]$ be the homomorphism of $\lambda$-rings as in [FSS24, Proposition 8.5.1(iii)].
@@
be the unique extension of the counting measure #\# in Theorem 3.3 sending $z$ to $z$. Similarly we abuse notation and let
@@
The homomorphism $ev_{X}$ factors through $\overline{\operatorname{Mot}}_{fin-type}^{\lambda-ring}(\mathbb{F}_{q})[[z]]$ and we have $\circ\mu=ev_{X}$.
@@
By [FSS24, Proposition 8.3.1(i)], $R_{g}$ is generated by the coefficients of $L^{univ}_{g}$, so it suffices to check the above diagram for $L^{univ}_{g}$ and $Q$. For $Q$ we have
@@
Note that the evaluation morphism $ev_{X}\colon K_{g}\to\mathbb{Q}$ sends the coefficient of $L^{univ}_{g}$ at $z^{i}$ to the coefficient of $L_{X}$ at $z^{i}$ and
$(L^{mot}_{X}(z))=L_{X}(z).$
By [FSS24, Proposition 8.5.1(i)], we have
$\mu\left(L^{univ}_{g}(z)\right)=L^{mot}_{X}(z)\in\overline{\operatorname{Mot}}_{fin-type}^{\lambda-ring}(\mathbb{F}_{q}).$
@@
Applying $\mu$ coefficient-wise to $\Omega_{g}^{univ}$, we get $\Omega_{X}^{mot}$, see [FSS24, Section 8.6]. Since $\mu\circ\operatorname{Log}_{g}=\operatorname{Log}\circ\mu$ and $\mu\circ\operatorname{Exp}_{g}=\operatorname{Exp}\circ\mu$, we have

### v2

@@ 5.1. Moduli stacks of vector bundles and HN-filtrations
    There is an open substack $\mathcal{B}un_{r,d}^{\geq\tau}\subset\mathcal{B}un_{r,d}$ classifying vector bundles whose HN-type ($\tau_{1}>\ldots>\tau_{t}$) satisfies $\tau_{t}\geq\tau$.
@@
    A vector bundle $E\in\mathcal{B}un_{r,d}(K)$ is in $\mathcal{B}un_{r,d}^{\geq\tau}(K)$ if and only if there is no surjective morphism of vector bundles $E\to F$ such that the slope of $F$ is less than $\tau$.
@@ 5.2. Moduli stacks of Higgs bundles
    A Higgs bundle on $X$ is a pair $(E,\Phi)$ where $E$ is a vector bundle on $X$ and $\Phi\colon E\to E\otimes\Omega_{X}$ is an $\mathcal{O}_{X}$-linear morphism from $E$ to $E$ “twisted” by the sheaf of differential $1$-forms $\Omega_{X}$.
@@
Similarly to the proof of [Lan75, Proposition 3] (see also Section 5.1), this is an open condition compatible with field extensions, so we denote by $\mathcal{H}iggs^{ss}_{r,d}$ the open substack of semistable Higgs bundles. Furthermore, the stack $\mathcal{H}iggs^{ss}_{r,d}$ is of finite type by [FSS18, Lemma 3.3.2(ii)].
@@ 5.3. Motivic Hall algebras
In this section, we keep the assumptions from the previous section, in particular: $X$ is a smooth projective geometrically connected curve of genus $g$ over $k$. In addition we assume that there is a divisor $D$ on $X$ defined over $k$ such that $\deg D=1$. In other words, the degrees of points on $X$ are coprime.
For any stack $\mathcal{X}$, we consider the ring
@@
Let $\mathcal{C}oh_{r,d}$ be the moduli stack of coherent sheaves on $X$ with rank $r$ and degree $d$. This is a stack locally of finite type over $k$ by [LMB00, Théorème 4.6.2.1]. Set
@@
Let us denote the group $\overline{\operatorname{Mot}}(\mathcal{X})$ defined in [FSS18, Section 2.1] without surjective and universally injective relation by $\overline{\operatorname{Mot}}^{naive}(\mathcal{X})$ and the corresponding motivic Hall algebra in [FSS18, Section 5.2] by $\mathcal{H}^{naive}$. Then we have the following well-defined homomorphisms
@@
Applying these homomorphisms to the identities in [FSS18, Section 5], we get identities in the “corrected” motivic Hall algebra such as [FSS18, Lemma 5.6.1 and Proposition 5.5.3]. The identity of [FSS18, Proposition 5.7.1] is also true in our $\mathcal{H}$: the argument goes through in arbitrary characteristic except for the part where the power structure is used, and this is where our Theorem 4.6 should be applied. After this, the calculations of [FSS18, Section 6] carry over to the case of finite characteristic. Thus for the remaining sections, we can use the results in loc. cit.
5.4. Motivic classes of the stacks of vector bundles with nilpotent endomorphisms
The following definitions come from [FSS18, Section 1.3] (note that we slightly re-write the definition of $J_{\lambda}^{mot}(z)$ as in [Mel20a, Section 3]). We define the “normalized” zeta-function $\tilde{\zeta}_{X}(z)\coloneqq z^{1-g}\zeta_{X}(z)$ where $g$ is the genus of the curve $X$.
Recall the definition of the motivic $L$-function in Section 1.6. Set
$J^{mot}_{\lambda}(z)\coloneqq\prod_{\square\in\lambda}\frac{L^{mot}_{X}(z^{a(\square)}\mathbb{L}^{-l(\square)-1})}{(1-z^{a(\square)}\mathbb{L}^{-l(\square)-1})(1-z^{a(\square)}\mathbb{L}^{-l(\square)})_{\neq 0}}\in\overline{\operatorname{Mot}}(k)[[z]],$
where the product is over all boxes of the Young diagram corresponding to the partition. Here the notation $(-)_{\neq 0}$ means that we omit this term if it is zero. In particular for the empty Young diagram $\lambda$ we have $J^{mot}_{\lambda}(z)=1$.
@@ 5.4. Motivic classes of the stacks of vector bundles with nilpotent endomorphisms
$L^{mot}(z_{n},\ldots,z_{1})\coloneqq\frac{1}{\prod_{i<j}\tilde{\zeta}_{X}\left(\frac{z_{i}}{z_{j}}\right)}\sum_{\sigma\in S_{n}}\sigma\left\{\prod_{i<j}\tilde{\zeta}_{X}\left(\frac{z_{i}}{z_{j}}\right)\frac{1}{\prod_{i<n}\left(1-\mathbb{L}\frac{z_{i+1}}{z_{i}}\right)}\cdot\frac{1}{1-z_{1}}\right\}.$
@@
Here the sum is over all partitions $\lambda=(\lambda_{1}\geq\lambda_{2}\geq\ldots\geq\lambda_{l}>0)$, $\langle\lambda,\lambda\rangle\coloneqq\sum_{i}(\lambda^{\prime}_{i})^{2}$ where $\lambda^{\prime}=(\lambda^{\prime}_{1}\geq\ldots\geq\lambda^{\prime}_{i}\geq\ldots)$ is the conjugate partition and $|\lambda|\coloneqq\sum_{i}\lambda_{i}$.
Let $\mathcal{E}nd_{r,d}^{nilp,+}$ denote the stack of HN-nonnegative bundles on $X$ of rank $r$ and degree $d$ with nilpotent endomorphisms.
@@
5.5. Explicit formulas for the motivic classes of the stacks of Higgs bundles
Finally we obtain the same formula for the motivic classes of the stacks of semistable Higgs bundles as in [FSS18, Theorem 1.3.3]. We remove the restriction on the characteristic of the field.
@@ 5.5. Explicit formulas for the motivic classes of the stacks of Higgs bundles
(5.2) $\sum_{r,d\in\mathbb{Z}_{\geq 0}\atop(r,d)\neq(0,0)}B_{r,d}w^{r}z^{d}=\mathbb{L}\operatorname{Log}\Omega_{X}^{Sch,mot}\in\overline{\operatorname{Mot}}(k)[[z,w]].$
@@
Recall that we have fixed a smooth projective geometrically connected curve $X$ of genus $g$. We have a result similar to [FSS18, Theorem 1.3.3] for the motivic classes of the moduli stacks of semistable Higgs bundles.
@@
Let $k$ be a field of arbitrary characteristic and let $X$ be a smooth projective geometrically connected curve over $k$. Assume that there exists a $k$-rational divisor of degree $1$ on $X$. We have
@@
whenever $d$ is large enough. This formula gives an explicit answer for the volumes $|\mathcal{H}iggs^{ss}_{r,d}(\mathbb{F}_{q})|$. It is not hard to derive this formula from our Theorems 3.2 and 5.5. We suppress the details because in the next section we will give a simpler formula for $|\mathcal{H}iggs^{ss}_{r,d}(\mathbb{F}_{q})|$ based on results of Mellit.
@@ 6. Mellit’s simplification
We would like to use Mellit’s results [Mel18, Mel20a] to obtain a formula for the motivic classes of the stacks of Higgs bundles in the universal $\lambda$-ring quotient that are much simpler than the formulas in Section 5.5. We also give a simpler formula for $|\mathcal{H}iggs^{ss}_{r,d}(\mathbb{F}_{q})|$.
6.1. Motivic classes of the stacks of Higgs bundles in the universal $\lambda$-ring quotient
Recall from Theorem 2.18 that we have a pre-$\lambda$-ring structure on the ring $\overline{\operatorname{Mot}}(k)$ such that for a reduced quasi-projective scheme $X$ we have
@@ 6.1. Motivic classes of the stacks of Higgs bundles in the universal $\lambda$-ring quotient
The following proposition is a no-puncture version of [FSS24, Corollary 8.7.4].
@@
Let $k$ be a field of arbitrary characteristic and let $X$ be a smooth projective geometrically connected curve over $k$. Assume that there exists a $k$-rational divisor of degree $1$ on $X$. Let $d/r=d_{0}/r_{0}$ where $r_{0}$ and $d_{0}$ are coprime integers with $r_{0}>0$. The image of the motivic class $\mathcal{H}iggs_{r,d}^{ss}]$ in $\overline{\operatorname{Mot}}^{\lambda-ring}(k)$ is equal to the coefficient at $w^{r}$ in
@@ 6.2. Counting Higgs bundles over finite fields
For convenience, we set $\alpha_{i+g}\coloneqq Q\alpha_{i}^{-1}$ for $i=1,\ldots,g$. We define a universal $L$-function
@@
$L_{X}(z)\coloneqq\left(\sum_{n\geq 0}|Sym^{n}X(\mathbb{F}_{q})|\cdot z^{n}\right)(1-z)(1-qz)$
is the “usual” $L$-function on $X$.
@@
Let $X$ be a smooth projective geometrically connected curve over $\mathbb{F}_{q}$. Assume that there exists an $\mathbb{F}_{q}$-rational divisor of degree $1$ on $X$. Let $d/r=d_{0}/r_{0}$ where $r_{0}$ and $d_{0}$ are coprime integers with $r_{0}>0$. The volume $|\mathcal{H}iggs^{ss}_{r,d}(\mathbb{F}_{q})|$ is the image under $ev_{X}$ of the coefficient at $w^{r}$ in
@@
Before giving the proof, we need some preliminaries. We define $V_{g}$ as the localization of $R_{g}[z]$ with respect to the multiplicative set generated by $Q^{i}-z^{j}$, where $i,j\geq 0$, $(i,j)\neq(0,0)$. Note that $K_{g}[z]\subset V_{g}$.
Let $\mu\colon V_{g}\to\overline{\operatorname{Mot}}^{\lambda-ring}(\mathbb{F}_{q})[[z]]$ be the homomorphism of $\lambda$-rings as in [FSS24, Proposition 8.5.1(iii)], sending $L^{univ}_{g}$ to $L^{mot}_{X}$ (see [FSS24, Proposition 8.5.1(i)]).
We denote by $\overline{\operatorname{Mot}}_{fin-type}^{\lambda-ring}(\mathbb{F}_{q})$ the image of the composition
By Theorem 3.2, the counting measure #\# factors through $\overline{\operatorname{Mot}}_{fin-type}^{\lambda-ring}(\mathbb{F}_{q})$.
@@
be the unique extension of #\# sending $z$ to $z$. Similarly we abuse notation and let
@@
We have $\circ\mu=ev_{X}$.
@@
By [FSS24, Proposition 8.3.1(i)], $R_{g}$ is generated by the coefficients of $L^{univ}_{g}$ as a $\mathbb{Z}[Q^{\pm 1}]$-algebra, so it suffices to check the commutativity of the above diagram for $L^{univ}_{g}$ and $Q$. For $Q$ we have
@@
Now we have the following sequence of equalities
$\circ\mu(L^{univ}_{g}(z))=\#(L^{mot}_{X}(z))=L_{X}(z)=ev_{X}(L^{univ}_{g}(z)),$
which follow from the definitions of $\mu$, of the counting measure #\#, and of $ev_{X}$ respectively. ∎
@@
Applying $\mu$ coefficient-wise to $\Omega_{g}^{univ}$, we get $\Omega_{X}^{mot}$, see [FSS24, Section 8.6]. Since $\mu$ is a homomorphism of pre-$\lambda$-rings, we have